OpenStudy (anonymous):

integral sqrt(9-4x^2)/x dx

OpenStudy (p0sitr0n):

integral sqrt(9-4 x^2)/x dx = sqrt(9-4 x^2)-3 log(sqrt(9-4 x^2)+3)+3 log(x)+c (constant)

OpenStudy (p0sitr0n):

if you need the steps, you can ask for them right away in the question. If you merely need the answer, you could check out wolframalpha.com

OpenStudy (solomonzelman):

u= 9-4x^2 du = -8x du you don't have the -8 there, so divide both sides by -8, you get -1/8 du= x dx

OpenStudy (anonymous):

Please, I need the steps

OpenStudy (solomonzelman):

So your substitution is u= 9-4x^2 -1/8 du = x dx

OpenStudy (solomonzelman):

I am telling you the steps no?

OpenStudy (solomonzelman):

I am not going to just do your work for you.

OpenStudy (solomonzelman):

I am not allowed to. But I can help you... if you ar willing to work.

OpenStudy (anonymous):

ok thanks

OpenStudy (solomonzelman):

Do you think you understand why I am substituting u=9-4x^2 and -1/8 du = x dx ?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

Have you learned U substitution?

OpenStudy (solomonzelman):

Lets do this.... \[\large \int\limits_{ }^{ } \frac{\sqrt{9-4x^2}}{x}~dx\]\[u=9-4x^2\]\[\large \frac{du}{dx}=0-2\times4x^{2-1}=-8x\] multiplying both sides times dx, \[du = -8x~ dx\]

OpenStudy (anonymous):

I think I did wrong exercise

OpenStudy (solomonzelman):

lets do this, so that you can understand the concept.

OpenStudy (solomonzelman):

or you don;t want to?

OpenStudy (solomonzelman):

I'll proceed doing this, and let me know of any changes.

OpenStudy (freckles):

I see a trig sub.

OpenStudy (solomonzelman):

ohh, it is 1/x there is no x inside.

OpenStudy (solomonzelman):

I am blind again

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

your u substitution gives a nice integrand for parital fractions

ganeshie8 (ganeshie8):

I prefer subing `9-4x^2 = u^2` though

OpenStudy (solomonzelman):

lol yes!!

OpenStudy (solomonzelman):

I haven't ever seen people do this, but don't see any restrictions against this. Awesome! You again taught me something valuable.

OpenStudy (solomonzelman):

then you would need to play around to get 1/x in term OF U.

ganeshie8 (ganeshie8):

you need an x in the numerator so you have to multiply x top and bottom

ganeshie8 (ganeshie8):

\[\large \int \frac{\sqrt{9-4x^2}}{x}~dx = \int \frac{\sqrt{9-4x^2}}{x^2}~xdx \]

OpenStudy (solomonzelman):

I don't understand what did you really do, and don't think arquimus gets it

OpenStudy (solomonzelman):

ohh, I see I thought of soemthing else... your square root didn't cover the entire thing, but that is the atex's fault.

OpenStudy (solomonzelman):

what now it does? lagggs ... hatwe it. Sorry for interrupting you,

OpenStudy (freckles):

he multiplied x on top and bottom

OpenStudy (solomonzelman):

yes, I got that part...

ganeshie8 (ganeshie8):

keep going with your substitution i think it goes smoothly, just chanege it to `9-4x^2 = u^2`

ganeshie8 (ganeshie8):

\[\large \int \frac{\sqrt{9-4x^2}}{x}~dx = \int \frac{\sqrt{9-4x^2}}{x^2}~xdx\] \(9-4x^2 = u^2 \implies xdx = \dfrac{-u~du}{4}\) \(x^2 = \frac{1}{4}(9-u^2)\) the integral becomes \[\large \int \dfrac{u^2}{u^2-9} ~du\]

OpenStudy (anonymous):

\[3\ln \left| \frac{ 3-\sqrt{9-4x^2} }{ x } \right| + \sqrt{9-4x^2 } + C\]

OpenStudy (anonymous):

I got this result, but I'm not okay

ganeshie8 (ganeshie8):

how did u get that ?

ganeshie8 (ganeshie8):

it looks correct except for a sign

OpenStudy (anonymous):

by trigonometric substitution

OpenStudy (anonymous):

if .... I think I made a mistake

OpenStudy (anonymous):

I will continue my efforts .... bye